# Cost-Sensitive Learning to Defer to Multiple Experts with Workload Constraints

### v2

@@ 5.2 Expert Decision Modeling - Performance and Calibration
Figure 2: Mean ECE and ROC-AUC for estimates of $\mathbb{P}(\mbox{y}_{i}=m_{j,i})$. Values are calculated for each expert $j$ and averaged, with error bars representing a 95(%) confidence interval. - Both methods obtain similar ROC-AUC, however, the value of $\lambda$ has significant impact on their ranking in terms of calibration, demonstrating the importance of testing L2D methods under a wide variety of cost-structures.
In the top row of Figure 2, we observe that the distribution of ROC-AUC is similar across both methods, with consistent overlap of error bars, despite a superior average across almost all scenarios for the HEM method. It is notable, however, that for both alert rates, the value of the ROC-AUC consistently falls below 0.50 when $\lambda=0.0114$, indicating that these models are not able to reliably rank probabilities of correctness of the expert team under these conditions. In these scenarios however, L2D methods can still theoretically achieve improvements in performance by choosing which instances should be deferred to the classifier $h$, whose probabilities of correctness have been shown to be reliably modeled. In the bottom row of Figure 2, we observe that HEM achieves significantly lower ECE for $\lambda=0.057$, with the reweighted OvA baseline outperforming HEM on scenarios with $\lambda=0.285$. This demonstrates how the performance ranking of different L2D methods can change significantly by altering the data distribution and cost-structure of the experimental setting.
@@ 5.3 Deferral under Capacity Constraints
In this section we evaluate the quality of deferral as measured by the average misclassification cost per 100 instances. The average experimental results are displayed in Table 3 with 95% confidence intervals, where we refer to our method as DeCCaF. We observe that, on average, both L2D methods outperform the non-L2D approaches, and that DeCCaF is significantly better in most scenarios. As the number of test seeds (25) is relatively low, the approximation of the 95% confidence interval may be unreliable, as the value shown is derived based on the central limit theorem.
Table 3: Expected Misclassification Cost per 100 instances ($\mathbb{E}[\mathcal{C}]/100$, assuming $c_{FP}=\lambda,c_{FN}=1$) for each {$a_{r}$,$\lambda$} pair. In each row, values are averaged across all 25 test variations, and displayed with 95% confidence intervals. FR and OC represent the “Full Rejection” and “Only Classifier” baselines.
@@
Table 4: % of the 25 Test Variations, for each {$a_{r}$,$\lambda$} pair, where DeCCaF beats other Methods. According to a binomial statistical significance test, with $\alpha=0.05$, DeCCaF is significantly better than other methods for values $\in[0.68,1]$, while values $\in[0.28,0.68]$ mean that we cannot conclude which method is best - Across all 6 scenarios, DeCCaF is shown to be superior in 5, while the comparison is inconclusive in 1.
@@
In Table 4, we calculate the percentage of test variations, for each scenario, in which our method outperforms other deferral baselines. These results demonstrate that, across all scenarios, DeCCaF outperforms the full rejection and the "Only Classifier" baseline. Notably, the relative performance of DeCCaF when compared to random assignment changes drastically depending on the scenario properties, performing no better than random assignment for $a_{r}=5\%FPR,\lambda=0.0114$. This again demonstrates how the cost structure and distribution of expert performances has significant impact on the performance of L2D systems. Nevertheless, we observe that DeCCaF performs significantly better than the OvA baseline in 5 out of 6 scenarios, outperforming in both scenarios with $\lambda=\lambda_{t}$. Note that the only scenario where the comparison between DeCCaF and other baselines is inconclusive corresponds to the most imbalanced scenario, where the prevalence is higher, and the cost of false positives is lowest. We therefore conclude that there are benefits in jointly modeling the human decision-making processes.
Figure 3: Expected Misclassification Cost per 100 instances ($\mathbb{E}[\mathcal{C}]/100$, assuming $c_{FP}=\lambda,c_{FN}=1$) for $a_{r}$ = {0.05,0.15},$\lambda=0.057$, and different amounts of training data. In each point, values are averaged across all 25 test variations, and displayed with 95% confidence intervals - DeCCaF remains significantly better in most scenarios.
@@
In Figure 3, we demonstrate the variations in misclassification cost as a result of the undersampling of training data. As expected, both methods are significantly impacted by reducing the amount of training data. We again observe that DeCCaF performs either similarly to, or outperforms the OvA baseline.
@@ 6 Conclusions and Future Work
In this work, we expand multi-expert L2D research to consider several real-world issues that limit adoption of these systems. We consider limited data availability, the existence of human expert work capacity constraints, cost-sensitive optimization objectives, and significant class imbalance, which are key challenges posed by many high-stakes classification tasks.
We propose a novel architecture, which aims to better model human behaviour, and to globally optimize the system’s performance under capacity constraints. We conducted constrained deferral under a wide variety of training and testing conditions, in a realistic cost-sensitive classification task, empirically demonstrating that variations in the cost structure, data distribution, and human behavior can have significant impact on the relative performance of deferral strategies. We demonstrate that DeCCaF performs significantly better than the baselines in a wide array of testing scenarios, showing that there are benefits to jointly modeling the expert team.
For future work, we plan on testing the addition of fairness incentives to our misclassification cost optimization method, to study the impact this system may have in ensuring fairness. For scenarios where the misclassification costs are instance-specific (i.e. transaction fraud), we will also study the possibility of direct cost estimation by using regression models instead of classification.
@@ A.1 Alert Model
As detailed in Section 4.3, our Alert Model is a LightGBM [Ke et al., 2017] classifier. The model was trained on the first 3 months of the BAF dataset, and validated on the fourth month. The model is trained by minimizing binary cross-entropy loss. The choice of hyperparameters is defined through Bayesian search [Akiba et al., 2019] on an extensive grid, for 100 trials, with validation done on the 4th month, where the optimization objective is to maximize recall at 5% false positive rate in validation. In Table  5 we present the hyperparameter search space used, as well as the parameters of the selected model.
@@ A.2 Determining $\lambda$ based on Neyman-Pearson criterion
In our task, the optimization goal is expressed by a Neyman-Pearson criterion, aiming to maximize recall subject to a fixed FPR of 5%. This criterion represents a trade-off between the cost of false positive $c_{FP}$ and the cost of false negative $c_{FN}$ errors. In bank account fraud prevention, the consequence of committing a false positive mistake, that is, rejecting a legitimate application, must be weighed against the cost of a false negative mistake, that is, accepting a fraudulent application. In Section 4.1, we state that our optimization goal is to obtain a set of predictions $\hat{y}$ that minimize the quantity
$\frac{1}{N}\sum_{i=1}^{N}\big{[}\lambda\mathbb{I}[\mbox{y}_{i}=0\land\hat{y}_{i}=1]+\mathbb{I}[\mbox{y}_{i}=1\land\hat{y}_{i}=0]\big{]}.$ (14)
However, in our task, we do not have access to the values of $c_{FP}$ and $c_{FN}$. To apply our misclassification cost re-weighting approach, we must then obtain the value $\lambda_{t}$ which is equivalent to the error cost trade-off enforced by the Neyman-Pearson criterion. This will allow us to set $c_{FP}=\lambda$ and $c_{FN}=1$.
According to Elkan [2001], we can establish a relationship between the ideal threshold of a binary classifier and the misclassification costs. For a given instance, the ideal classification is the one that minimizes the expected loss. As such, the optimal prediction for any given instance $\bm{x_{i}}$ is 1 only if the expected cost of predicting 1 is less than, or equal to the expected cost of predicting 0, which is equivalent to:
$(1-p)c_{FP}\leq pc_{FN}$ (15)
@@
$(1-t)c_{FP}=tc_{FN}\Leftrightarrow\lambda_{t}=\frac{t}{1-t}=0.057$ (16)
As the optimal threshold $t$ for our Alert Model was chosen such that the Neyman-Pearson criterion is met, we now may plug the value of $t$ into this equation, obtaining the theoretical value $\lambda_{t}$ for our optimization goal. It has been shown by Sheng and Ling [2006] that this relationship between the cost-structure and the classifier’s threshold does not always hold in practice. Secondly, the value of $\lambda$ obtained through this method depends on the classifier trained. A different classifier would yield another value for the optimal threshold according to the Neyman-Pearson criterion, which would lead to a different $\lambda$, despite the task being the same. However, as our aim is to test different cost structures, we will use this as the default value for $\lambda$, testing the cost-structures $\lambda\in\{\lambda_{t}/5,\lambda_{t},5\lambda_{t}\}$
@@ Appendix B Classifier h
As detailed in Section 4.1, we model the classifier $h$ with a LightGBM [Ke et al., 2017] classifier. The model is trained on the alerts raised by the Alert Model, ranging from months four to six of the BAF dataset. The model is trained by minimizing the weighted log-loss as mentioned in section 3.1. The choice of hyperparameters is defined through Bayesian search [Akiba et al., 2019] on an extensive grid, for 300 total trials, with 200 start-up trials, with validation done on the seventh month. We also test several values for the initial probability estimate of the base predictor of the boosting model. This was done in order to test if introducing a bias towards predicting fraud can be beneficial to our model, as across all scenarios, false negatives incur a higher cost. Given instance-wise feature weights $c_{i}$, the default value $g_{\mbox{\scriptsize initial}}$ of the initial estimator’s prediction is
$g_{d}=logit\bigg{(}\frac{\sum_{i}c_{i}\mbox{y}_{i}}{\sum_{i}\mbox{y}_{i}}\bigg{)}.$ (17)
@@ C.1 Expert Desiderata
However, in some real world deferral systems [Amarasinghe et al., 2022, De-Arteaga et al., 2020], the instance’s features are accompanied by an AI model’s score, representing the model’s estimate of the probability that said instance belongs to the positive class. The aim of presenting the model score to an expert is to provide them with extra information, as well as possibly expediting the decision process. It has been shown that, in this scenario, expert’s performance can be impacted by presenting the model’s score when deferring a case to an expert [Amarasinghe et al., 2022, De-Arteaga et al., 2020, Levy et al., 2021]. Note that, in this setting, a classifier trained for our task exists independently of the assignment system implemented. This approach is applicable to deferral systems such as the one proposed by Raghu et al. [2019a], who argue for the training of a separate human behaviour prediction model [Raghu et al., 2019a, b]. In the L2D framework, however, the main classifier is trained jointly with the deferral system. As such, this framework allows users to generate experts with or without dependence on a separate ML classifier’s score.
@@
Should the generated experts use an AI assistant, we expect experts not to be in perfect agreement with the model, due to the assumption that humans and models have complementary strengths and weaknesses [De-Arteaga et al., 2020, Dellermann et al., 2019]. As such, we would assume humans and AI perform better than one another in separate regions of the feature space, enabling an assignment system to obtain better performance than either the expert team or the model on their own. The degree of "model dependence", or "algorithmic bias" [Alon-Barkat and Busuioc, 2023],varies between humans, as measured by the model’s impact on a human’s performance [Jacobs et al., 2021, Inkpen et al., 2022]. As such, our synthetic expert team may also exhibit varying levels of dependence on the model score.
@@
In order for our team of experts to be realistic, it is important that these exhibit varying levels of overall performance. Experts within a field have been shown to have varying degrees of expertise, with some being outperformed by ML models [Goel et al., 2021, Gulshan et al., 2016]. As such human decision processes can be expected to be varied even amongst a team of experts.
@@
It is a common assumption that, when making a decision, experts follow an internal process based on the available information. However, it is also known that highly educated and trained individuals are still subject to flaws that are inherent to human decision making processes, one of these being inconsistency. When presented with similar cases, at different times, experts may perform drastically different decisions [Danziger et al., 2011, Grimstad and Jørgensen, 2007]. Therefore we can expect a human’s decision making process not to be entirely deterministic.
@@ C.2 Expert Parameter Sampling
$\begin{gathered}\begin{cases}\mathbb{P}(m_{j,i}=1\lvert\mbox{y}_{i}=0,\bm{x}_{i})=\sigma\Big{(}\beta_{0}-\alpha\frac{\bm{w}\cdot\bm{\bar{x}}_{i}+w_{M}M(\bm{x}_{i})}{\sqrt{\bm{w}\cdot\bm{w}+w_{M}^{2}}}\Big{)}\\ \mathbb{P}(m_{j,i}=0\lvert\mbox{y}_{i}=1,\bm{x}_{i})=\sigma\Big{(}\beta_{1}+\alpha\frac{\bm{w}\cdot\bm{\bar{x}}_{i}+\bm{w}_{M}M(\bm{x}_{i})}{\sqrt{\bm{w}\cdot\bm{w}+w_{M}^{2}}}\Big{)},\end{cases}\end{gathered}$ (18)
@@
To define $\bm{w}$ for a given expert, we sample each component from a "Spike and Slab" prior [Mitchell and Beauchamp, 1988]. A spike and slab prior is a generative model in which a random variable $u$ either attains some fixed value $v$, called the spike, or is drawn from another prior $p_{\mbox{\scriptsize slab}}$, called the slab. In our case, we set $v=0$. That is, $u$ is either zero, or drawn from the slab density $N(0,1)$. To sample the values of $w_{i}$, we first sample a Bernoulli latent variable $Z\sim Ber(0.3)$ to select if $w_{i}$ is sampled from the spike or the slab. If $Z=0$, $w_{i}$ attains the fixed value $v=0$, if $Z=1$, $w_{i}$ is drawn from the slab density $p_{\mbox{\scriptsize slab}}$. As such, the spike and slab prior induces sparsity unless $\theta=1$, allowing for the generation of experts whose probabilities of error are swayed by a varying number of features. The distribution of $w_{M}\sim N(-2,0.5)$ is defined separately to ensure all experts have some degree of model dependency. We also separately define $w_{p}\sim N(-1,0.1)$, with $p$ representing the protected attribute, that is, the customer’s age, so that all experts have a degree of unfairness in their simulated predictions, however, an exploration of fairness was not conducted in this work.
@@
$\begin{split}\mathbb{E}[\mathcal{C}]_{j}\approx\frac{1}{N}\sum_{i=1}^{N}&\big{[}\lambda\mathbb{P}(m_{j,i}=1\lvert\mbox{y}_{i}=0)\mathbb{P}(\mbox{y}_{i}=0)\\ &+\mathbb{P}(m_{j,i}=0\lvert\mbox{y}_{i}=1)\mathbb{P}(\mbox{y}_{i}=1)\big{]},\\ \end{split}$ (19)
where $\mathbb{P}(m_{j,i}=1\lvert\mbox{y}_{i}=0)$ and $\mathbb{P}(m_{j,i}=0\lvert\mbox{y}_{i}=1)$ are given by Equation 18. Assuming that the values of $\bm{w},\alpha$ and $w_{M}$ are set, the value of $\mathbb{E}[C]_{j}$ will only be a function of $\beta_{0}$ and $\beta_{1}$. We must then be able to calculate values of $\beta_{0}$ and $\beta_{1}$ so that we obtain $\mathbb{E}[C]_{j}=T_{\mathbb{E}[C]_{j}}$. Let us simplify the above equation
$\begin{split}T_{\mathbb{E}[C]_{j}}&=\lambda\mathbb{P}(\mbox{y}_{i}=0)\frac{1}{N}\sum_{i=1}^{N}\big{[}\mathbb{P}(m_{j,i}=1\lvert\mbox{y}_{i}=0)\big{]}\\ &+\mathbb{P}(\mbox{y}_{i}=1)\frac{1}{N}\sum_{i=1}^{N}\big{[}\mathbb{P}(m_{j,i}=0\lvert\mbox{y}_{i}=1)\big{]},\\ &=\lambda(1-\mathbb{P}(\mbox{y}_{i}=1))\;\mbox{FPR}_{j}+\mathbb{P}(\mbox{y}_{i}=1)\;\mbox{FNR}_{j},\end{split}$ (20)
@@
$\begin{split}\mbox{FPR}_{j}=\frac{T_{\mathbb{E}[C]_{j}}}{\lambda(1-\mathbb{P}(\mbox{y}_{i}=1))}-\frac{\mathbb{P}(\mbox{y}_{i}=1)}{\lambda(1-\mathbb{P}(\mbox{y}_{i}=1))}\mbox{FNR}_{j},\end{split}$ (21)
@@
$\frac{\partial\mbox{FPR}_{j}}{\partial\beta_{1}}=\frac{1}{N}\sum_{i}\sigma\Big{(}\beta_{1}+\alpha\frac{\bm{w}.\bm{x}_{i}}{\lvert\lvert\bm{w}\lvert\lvert}\Big{)}\Bigg{(}1-\sigma\Big{(}\beta_{1}+\alpha\frac{\bm{w}.\bm{x}_{i}}{\lvert\lvert\bm{w}\lvert\lvert}\Big{)}\Bigg{)}>0\quad\mbox{for}\quad\beta\in\mathbb{R}\;.$ (23)
@@
$\mbox{FPR}_{j}(\beta_{1})-T_{\mbox{\scriptsize FPR}}=0\quad\mbox{for}\quad T_{\mbox{\scriptsize FPR}}\in\;]0,1[\;.$ (24)
A similar reasoning applies for $\beta_{0}$ and $T_{\mbox{ \scriptsize FNR}}$. Finally, we can control an expert’s FPR and FNR by solving these equations for $\beta_{1}$ and $\beta_{0}$. To solve these equations, a partition of the dataset (month 7) is utilized to calculate the empirical value for each rate, meaning that deviations from the target performance metrics may occur due to temporal distribution shifts. Due to the monotonous nature of the function, and the uniqueness of the solution, we solve it through a bisection method [Burden and Faires, 1985].

### v3

@@ 5.2 Expert Decision Modeling - Performance and Calibration
Figure 2: Mean ECE and ROC-AUC for estimates of $\mathbb{P}(y_{i}=m_{j,i})$. Values are calculated for each expert $j$ and averaged, with error bars representing a 95(%) confidence interval. Both methods obtain similar ROC-AUC, however, the value of $\lambda$ has significant impact on their ranking in terms of calibration, showing the importance of testing L2D methods under a wide variety of cost-structures.
In the top row of Figure 2, we observe that the distribution of ROC-AUC is similar across both methods, with consistent overlap of error bars, despite a superior average across almost all scenarios for the HEM method. It is notable, however, that for both alert rates, the value of the ROC-AUC consistently falls below 0.50 when $\lambda=0.0114$, indicating that these models are not able to reliably rank probabilities of correctness of the expert team under these conditions. In these scenarios, however, L2D methods can still theoretically achieve improvements in performance by choosing which instances should be deferred to the classifier $h$, whose probabilities of correctness have been shown to be reliably modeled. In the bottom row of Figure 2, we observe that HEM achieves significantly lower ECE for $\lambda=0.057$, with the reweighted OvA baseline outperforming HEM in scenarios with $\lambda=0.285$. This demonstrates how the performance ranking of different L2D methods can change significantly by altering the data distribution and cost-structure of the experimental setting.
@@ 5.3 Deferral under Capacity Constraints
In this section, we evaluate the quality of deferral as measured by the average misclassification cost per 100 instances. The average experimental results are displayed in Table 3 with 95% confidence intervals, where we refer to our method as DeCCaF. We observe that, on average, both L2D methods outperform the non-L2D approaches, and that DeCCaF is significantly better in most scenarios. As the number of test seeds (25) is relatively low, the approximation of the 95% confidence interval may be unreliable, as the value shown is derived based on the central limit theorem.
Table 3: Expected misclassification cost per 100 instances ($\mathbb{E}[\mathcal{C}]/100$, assuming $c_{\mbox{\scriptsize FP}}=\lambda,c_{\mbox{\scriptsize FN}}=1$) for each {$a_{r}$,$\lambda$} pair. In each row, values are averaged across all 25 test variations, and displayed with 95% confidence intervals. FR and OC represent the “Full Rejection” and “Only Classifier” baselines.
@@
Table 4: Percentage of the 25 test variations, for each {$a_{r}$,$\lambda$} pair, where DeCCaF beats other methods. According to a binomial statistical significance test, with $\alpha=0.05$, DeCCaF is significantly better than other methods for values $\in[0.68,1]$, while values $\in[0.28,0.68]$ mean that we cannot conclude which method is best. Across all 6 scenarios, DeCCaF is shown to be superior in 5, while the comparison is inconclusive in one.
@@
In Table 4, we report the percentage of test variations, for each scenario, in which our method outperforms other deferral baselines. These results demonstrate that, across all scenarios, DeCCaF outperforms the full rejection and the "Only Classifier" baseline. Notably, the relative performance of DeCCaF when compared to random assignment changes drastically depending on the scenario properties, performing no better than random assignment for $a_{r}=5\%FPR,\lambda=0.0114$. This again demonstrates how the cost structure and distribution of expert performances has significant impact on the performance of L2D systems. Nevertheless, we observe that DeCCaF performs significantly better than the OvA baseline in 5 out of 6 scenarios, outperforming in both scenarios with $\lambda=\lambda_{t}$. Note that the only scenario where the comparison between DeCCaF and other baselines is inconclusive corresponds to the most imbalanced scenario, where the prevalence is higher, and the cost of false positives is lowest. We therefore conclude that there are benefits in jointly modeling the human decision-making processes.
@@
Figure 3: Expected Misclassification Cost per 100 instances ($\mathbb{E}[\mathcal{C}]/100$, assuming $c_{\mbox{\scriptsize FP}}=\lambda,c_{\mbox{\scriptsize FN}}=1$) for $a_{r}\in\{0.05,0.15\}$, $\lambda=0.057$, and different amounts of training data. In each point, values are averaged across all 25 test variations, and displayed with 95% confidence intervals - DeCCaF remains significantly better in most scenarios.
In Figure 3, we show the variations in misclassification cost as a result of the undersampling of training data. As expected, both methods are significantly impacted by reducing the amount of training data. We again observe that DeCCaF performs either similarly to, or outperforms the OvA baseline.
@@ 6 Conclusions and Future Work
In this work, we expand multi-expert L2D research to consider several real-world issues that limit adoption of these systems. We consider limited data availability, the existence of human expert work-capacity constraints, cost-sensitive optimization objectives, and significant class imbalance, which are key challenges posed by many high-stakes classification tasks.
We propose a novel architecture that aims to better model human behaviour and to globally optimize the system’s performance under capacity constraints. We conducted constrained deferral under a wide variety of training and testing conditions, in a realistic cost-sensitive classification task, empirically demonstrating that variations in the cost structure, data distribution, and human behavior can have significant impact on the relative performance of deferral strategies. We demonstrate that DeCCaF performs significantly better than the baselines in a wide array of testing scenarios, showing that there are benefits to jointly modeling the expert team.
For future work, we plan on testing the addition of fairness incentives to our misclassification cost optimization method, to study the impact this system may have in ensuring fairness. For scenarios where the misclassification costs are instance-specific (i.e., transaction fraud), we will also study the possibility of direct cost estimation by using regression models instead of classification.
@@ A.1 Alert Model
As detailed in Section 4.3, our Alert Model is a LightGBM (Ke et al., 2017) classifier. The model was trained on the first 3 months of the BAF dataset, and validated on the fourth month. The model is trained by minimizing binary cross-entropy loss. The choice of hyperparameters is defined through Bayesian search (Akiba et al., 2019) on an extensive grid, for 100 trials, with validation done on the 4th month, where the optimization objective is to maximize recall at 5% false positive rate in validation. In Table  5 we present the hyperparameter search space used, as well as the parameters of the selected model.
@@ A.2 Determining $\lambda$ based on Neyman-Pearson criterion
In our task, the optimization goal is expressed by a Neyman-Pearson criterion, aiming to maximize recall subject to a fixed FPR of 5%. This criterion represents a trade-off between the cost of false positive $c_{\mbox{\scriptsize FP}}$ and the cost of false negative $c_{\mbox{\scriptsize FN}}$ errors. In bank account fraud prevention, the consequence of committing a false positive mistake, that is, rejecting a legitimate application, must be weighed against the cost of a false negative mistake, that is, accepting a fraudulent application. In Section 4.1, we state that our optimization goal is to obtain a set of predictions $\hat{y}$ that minimize the quantity
$\frac{1}{N}\sum_{i=1}^{N}\big{[}\lambda\mathbb{I}[y_{i}=0\land\hat{y}_{i}=1]+\mathbb{I}[y_{i}=1\land\hat{y}_{i}=0]\big{]}.$ (19)
However, in our task, we do not have access to the values of $c_{\mbox{\scriptsize FP}}$ and $c_{\mbox{\scriptsize FN}}$. To apply our misclassification cost re-weighting approach, we must then obtain the value $\lambda_{t}$ which is equivalent to the error cost trade-off enforced by the Neyman-Pearson criterion. This will allow us to set $c_{\mbox{\scriptsize FP}}=\lambda$ and $c_{\mbox{\scriptsize FN}}=1$.
According to Elkan (2001), we can establish a relationship between the ideal threshold of a binary classifier and the misclassification costs. For a given instance, the ideal classification is the one that minimizes the expected loss. As such, the optimal prediction for any given instance $\bm{x_{i}}$ is 1 only if the expected cost of predicting 1 is less than, or equal to the expected cost of predicting 0, which is equivalent to:
$(1-p)c_{\mbox{\scriptsize FP}}\leq pc_{\mbox{\scriptsize FN}}$ (20)
@@
$(1-t)c_{\mbox{\scriptsize FP}}=tc_{\mbox{\scriptsize FN}}\Leftrightarrow\lambda_{t}=\frac{t}{1-t}=0.057$ (21)
As the optimal threshold $t$ for our Alert Model was chosen such that the Neyman-Pearson criterion is met, we now may plug the value of $t$ into this equation, obtaining the theoretical value $\lambda_{t}$ for our optimization goal. It has been shown by Sheng & Ling (2006) that this relationship between the cost-structure and the classifier’s threshold does not always hold in practice. Secondly, the value of $\lambda$ obtained through this method depends on the classifier trained. A different classifier would yield another value for the optimal threshold according to the Neyman-Pearson criterion, which would lead to a different $\lambda$, despite the task being the same. However, as our aim is to test different cost structures, we will use this as the default value for $\lambda$, testing the cost-structures $\lambda\in\{\lambda_{t}/5,\lambda_{t},5\lambda_{t}\}$
@@ Appendix B Classifier h
As detailed in Section 4.1, we model the classifier $h$ with a LightGBM (Ke et al., 2017) classifier. The model is trained on the alerts raised by the Alert Model, ranging from months four to six of the BAF dataset. The model is trained by minimizing the weighted log-loss as mentioned in section 3.1. The choice of hyperparameters is defined through Bayesian search (Akiba et al., 2019) on an extensive grid, for 300 total trials, with 200 start-up trials, with validation done on the seventh month. We also test several values for the initial probability estimate of the base predictor of the boosting model. This was done in order to test if introducing a bias towards predicting fraud can be beneficial to our model, as across all scenarios, false negatives incur a higher cost. Given instance-wise feature weights $c_{i}$, the default value $g_{\mbox{\scriptsize initial}}$ of the initial estimator’s prediction is
$g_{d}=logit\bigg{(}\frac{\sum_{i}c_{i}y_{i}}{\sum_{i}y_{i}}\bigg{)}.$ (22)
@@ C.1 Expert Desiderata
However, in some real world deferral systems (Amarasinghe et al., 2022; De-Arteaga et al., 2020), the instance’s features are accompanied by an AI model’s score, representing the model’s estimate of the probability that said instance belongs to the positive class. The aim of presenting the model score to an expert is to provide them with extra information, as well as possibly expediting the decision process. It has been shown that, in this scenario, expert’s performance can be impacted by presenting the model’s score when deferring a case to an expert (Amarasinghe et al., 2022; De-Arteaga et al., 2020; Levy et al., 2021). Note that, in this setting, a classifier trained for our task exists independently of the assignment system implemented. This approach is applicable to deferral systems such as the one proposed by Raghu et al. (2019a), who argue for the training of a separate human behaviour prediction model (Raghu et al., 2019a; b). In the L2D framework, however, the main classifier is trained jointly with the deferral system. As such, this framework allows users to generate experts with or without dependence on a separate ML classifier’s score.
@@
Should the generated experts use an AI assistant, we expect experts not to be in perfect agreement with the model, due to the assumption that humans and models have complementary strengths and weaknesses (De-Arteaga et al., 2020; Dellermann et al., 2019). As such, we would assume humans and AI perform better than one another in separate regions of the feature space, enabling an assignment system to obtain better performance than either the expert team or the model on their own. The degree of "model dependence", or "algorithmic bias" (Alon-Barkat & Busuioc, 2023),varies between humans, as measured by the model’s impact on a human’s performance (Jacobs et al., 2021; Inkpen et al., 2022). As such, our synthetic expert team may also exhibit varying levels of dependence on the model score.
@@
In order for our team of experts to be realistic, it is important that these exhibit varying levels of overall performance. Experts within a field have been shown to have varying degrees of expertise, with some being outperformed by ML models (Goel et al., 2021; Gulshan et al., 2016). As such human decision processes can be expected to be varied even amongst a team of experts.
@@
It is a common assumption that, when making a decision, experts follow an internal process based on the available information. However, it is also known that highly educated and trained individuals are still subject to flaws that are inherent to human decision making processes, one of these being inconsistency. When presented with similar cases, at different times, experts may perform drastically different decisions (Danziger et al., 2011; Grimstad & Jørgensen, 2007). Therefore we can expect a human’s decision making process not to be entirely deterministic.
@@ C.2 Expert Parameter Sampling
$\begin{gathered}\begin{cases}\mathbb{P}(m_{j,i}=1\lvert y_{i}=0,\bm{x}_{i})=\sigma\Big{(}\beta_{0}-\alpha\frac{\bm{w}\cdot\bm{\bar{x}}_{i}+w_{M}M(\bm{x}_{i})}{\sqrt{\bm{w}\cdot\bm{w}+w_{M}^{2}}}\Big{)}\\ \mathbb{P}(m_{j,i}=0\lvert y_{i}=1,\bm{x}_{i})=\sigma\Big{(}\beta_{1}+\alpha\frac{\bm{w}\cdot\bm{\bar{x}}_{i}+\bm{w}_{M}M(\bm{x}_{i})}{\sqrt{\bm{w}\cdot\bm{w}+w_{M}^{2}}}\Big{)},\end{cases}\end{gathered}$ (23)
@@
To define $\bm{w}$ for a given expert, we sample each component from a "Spike and Slab" prior (Mitchell & Beauchamp, 1988). A spike and slab prior is a generative model in which a random variable $u$ either attains some fixed value $v$, called the spike, or is drawn from another prior $p_{\mbox{\scriptsize slab}}$, called the slab. In our case, we set $v=0$. That is, $u$ is either zero, or drawn from the slab density $N(0,1)$. To sample the values of $w_{i}$, we first sample a Bernoulli latent variable $Z\sim Ber(0.3)$ to select if $w_{i}$ is sampled from the spike or the slab. If $Z=0$, $w_{i}$ attains the fixed value $v=0$, if $Z=1$, $w_{i}$ is drawn from the slab density $p_{\mbox{\scriptsize slab}}$. As such, the spike and slab prior induces sparsity unless $\theta=1$, allowing for the generation of experts whose probabilities of error are swayed by a varying number of features. The distribution of $w_{M}\sim N(-2,0.5)$ is defined separately to ensure all experts have some degree of model dependency. We also separately define $w_{p}\sim N(-1,0.1)$, with $p$ representing the protected attribute, that is, the customer’s age, so that all experts have a degree of unfairness in their simulated predictions, however, an exploration of fairness was not conducted in this work.
@@
$\begin{split}\mathbb{E}[\mathcal{C}]_{j}\approx\frac{1}{N}\sum_{i=1}^{N}&\big{[}\lambda\mathbb{P}(m_{j,i}=1\lvert y_{i}=0)\mathbb{P}(y_{i}=0)\\ &+\mathbb{P}(m_{j,i}=0\lvert y_{i}=1)\mathbb{P}(y_{i}=1)\big{]},\\ \end{split}$ (24)
where $\mathbb{P}(m_{j,i}=1\lvert y_{i}=0)$ and $\mathbb{P}(m_{j,i}=0\lvert y_{i}=1)$ are given by Equation 23. Assuming that the values of $\bm{w},\alpha$ and $w_{M}$ are set, the value of $\mathbb{E}[C]_{j}$ will only be a function of $\beta_{0}$ and $\beta_{1}$. We must then be able to calculate values of $\beta_{0}$ and $\beta_{1}$ so that we obtain $\mathbb{E}[C]_{j}=T_{\mathbb{E}[C]_{j}}$. Let us simplify the above equation
$\begin{split}T_{\mathbb{E}[C]_{j}}&=\lambda\mathbb{P}(y_{i}=0)\frac{1}{N}\sum_{i=1}^{N}\big{[}\mathbb{P}(m_{j,i}=1\lvert y_{i}=0)\big{]}\\ &+\mathbb{P}(y_{i}=1)\frac{1}{N}\sum_{i=1}^{N}\big{[}\mathbb{P}(m_{j,i}=0\lvert y_{i}=1)\big{]},\\ &=\lambda(1-\mathbb{P}(y_{i}=1))\;\mbox{FPR}_{j}+\mathbb{P}(y_{i}=1)\;\mbox{FNR}_{j},\end{split}$ (25)
@@
$\begin{split}\mbox{FPR}_{j}=\frac{T_{\mathbb{E}[C]_{j}}}{\lambda(1-\mathbb{P}(y_{i}=1))}-\frac{\mathbb{P}(y_{i}=1)}{\lambda(1-\mathbb{P}(y_{i}=1))}\mbox{FNR}_{j},\end{split}$ (26)
@@
$\frac{\partial\mbox{FPR}_{j}}{\partial\beta_{1}}=\frac{1}{N}\sum_{i}\sigma\Big{(}\beta_{1}+\alpha\frac{\bm{w}.\bm{x}_{i}}{\lvert\lvert\bm{w}\lvert\lvert}\Big{)}\Bigg{(}1-\sigma\Big{(}\beta_{1}+\alpha\frac{\bm{w}.\bm{x}_{i}}{\lvert\lvert\bm{w}\lvert\lvert}\Big{)}\Bigg{)}>0\;\mbox{for}\;\beta\in\mathbb{R}\;.$ (28)
@@
$\mbox{FPR}_{j}(\beta_{1})-T_{\mbox{\scriptsize FPR}}=0\;\mbox{for}\;T_{\mbox{\scriptsize FPR}}\in\;]0,1[\;.$ (29)
A similar reasoning applies for $\beta_{0}$ and $T_{\mbox{ \scriptsize FNR}}$. Finally, we can control an expert’s FPR and FNR by solving these equations for $\beta_{1}$ and $\beta_{0}$. To solve these equations, a partition of the dataset (month 7) is utilized to calculate the empirical value for each rate, meaning that deviations from the target performance metrics may occur due to temporal distribution shifts. Due to the monotonous nature of the function, and the uniqueness of the solution, we solve it through a bisection method (Burden & Faires, 1985).